OpenStudy (s117):

. solve for x A. –14 B. –10 C. –3 D. 1

OpenStudy (s117):

\[\frac{ x+2 }{ 3 }=-4\]

OpenStudy (anonymous):

First we multiply 3 to both sides, what's -4 * 3?

OpenStudy (anonymous):

^ @S117

OpenStudy (s117):

srry was away

OpenStudy (anonymous):

It's okay, multiply -4 * 3

OpenStudy (s117):

-12

OpenStudy (anonymous):

Yes, that gives us x + 2 = -12 Now subtract 2 to both sides, what's -12 - 2?

OpenStudy (s117):

+2 so -10 right?

OpenStudy (anonymous):

No, not exactly.

OpenStudy (s117):

its that or -14

OpenStudy (anonymous):



OpenStudy (anonymous):

-14 is correct.

OpenStudy (s117):

yay